# Efficient and Effective Query Context-Aware Learning-to-Rank Model for Sequential Recommendation

### v1

@@ Abstract.
Modern sequential recommender systems commonly use transformer-based models for next-item prediction. While these models demonstrate a strong balance between efficiency and quality, integrating interleaving features – such as the query context (e.g., browse category) under which next-item interactions occur – poses challenges. Effectively capturing query context is crucial for refining ranking relevance and enhancing user engagement, as it provides valuable signals about user intent within a session. Unlike an item’s features, query context is not temporally aligned with the item sequence, making its incorporation into transformers challenging and error-prone. This paper analyzes different strategies for incorporating query context into transformers trained with a causal language modeling procedure as a case study. We propose a new method that effectively fuses the item sequence with query context within the attention mechanism. Through extensive offline and online experiments on a large-scale online platform and open datasets, we present evidence that our proposed method is an effective approach for integrating query context to improve model ranking quality in terms of relevance and diversity.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Nineteenths ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicccs: Information systems Recommender systems
@@ 1. Introduction
Existing research on efficient transformer-based sequential models primarily focuses on integrating an item’s features as contextual information (Yang et al., 2021; Cai et al., 2021) or treating query context as side information in conjunction with the user representation, without fusing it early with item sequence (Covington et al., 2016a; Zhou et al., 2018), which may limit model effectiveness. When historical query context is assumed to be available online, it has been effectively incorporated into recurrent neural network-based approaches by embedding it within the item sequence, thereby fusing it with the user representation ((Liu et al., 2016; Song et al., 2017; Beutel et al., 2018)).
In this paper, we propose integrating query context into sequential recommendation models in an efficient and effective manner. Specifically, we address the following research questions:
@@
To address these questions, we explore various methods for incorporating query context into transformer-based sequential recommenders, using causal language modeling as a case study.
Our key contributions are as follows:
  • A comprehensive study on query context integration, focusing on both efficiency and effectiveness.
  • A novel approach for incorporating query context into transformer-based ranking models. Our method enhances the expressiveness of the user representation by contextually fusing it with query context, leading to improved model accuracy.
  • Extensive offline and online experiments on a large-scale online platform. We validate our findings using open datasets and provide open-source code for reproducibility.
@@ 2. Related work
Query context is typically incorporated as side information alongside the user representation, rather than being fused early with the item sequence (Covington et al., 2016a; Zhou et al., 2018). However, treating query context as side information may limit the predictive power of the user representation, as it is not explicitly tailored to the query context.
@@ 3.1. Problem statement
During estimation, the $(N+1)$th item is used only as the target when predicting the $N$th next item, while the first item prediction is used for cold-start users.
@@ 3.2. Method
To efficiently model the problem, we consider the deep self-attention transformer model trained with causal language modeling (Kang and McAuley, 2018; Li et al., 2020; Zhang et al., 2019). A transformer begins with an item embedding, followed by a series of self-attention layers, and concludes with an item unembedding. These models stack multiple self-attention layers, each followed by a position-wise feed-forward sublayer, residual connections, and layer normalization (Ba et al., 2016; Nguyen and Salazar, 2019). Multi-head self-attention is commonly used to capture different aspects of input sequences (Vaswani et al., 2017), though we omit it here for notational simplicity.
@@ 4.1. Fixing temporal misalignment
where $\bm{L}$ is a modified lower shift binary $N$-by-$(N+1)$ matrix with ones on the subdiagonal. The projections (1) in subsequent layers $h=2,\dots,H$ remain unchanged. This ensures $c_{i+1}$ is correctly placed at query position $\bm{Q}^{(1)}_{i}$, influencing attention over past items $\left(x_{t}^{u}\right)_{t=1}^{i}$ while also propagating the query context further through keys and values. However, this approach raises concerns:
@@ 5.3. Approach C: Query context in the last layer’s query position and outside
Analyzing attention flow in causal language modeling, we conclude that the safest way to introduce query context within attention without causing unintended misrepresentations (Section 4.1) is by placing it in the last layer, exclusively at the query position.
@@
while the projections in the previous layers, $11=2,\dots,H-$ 1, remain unchanged as defined in Eq. (1).
@@ 6. Experiments on our large-scale online platform
The model follows a two-tower architecture (Covington et al., 2016b; Yang et al., 2020), with the training procedure described in Section 3.2. It consists of a user tower, represented by a transformer with residual layer blocks, and an item tower, represented by a feed-forward network, whose output is the item embedding $\bm{A}$ from Eq. 3. Although these towers are trained jointly, they are deployed and operated independently. The item embeddings are indexed in a vector store for efficient ranking and retrieval. Each time a user makes a request, the user tower generates a new user embedding based on their most recent interactions (up to the last 100 items). This embedding is then used to retrieve the most relevant items using a nearest-neighbor index.
@@ 6.2. Online experiments
    For high-value actions (add-to-wishlist and add-to-cart), both new and returning users saw positive impacts: +1.0% [0.2, 1.8]for new and +1.7% [1.4 , 2.0]for returning users. This suggests that the treatment benefits even users for whom we have little to no prior data.
@@ 8. Discussion
We propose three approaches: Approach B, which partially mitigates the mismatch through query context masking, and Approaches A and C, which inherently avoid the training-serving mismatch.

### v2

@@ Abstract.
Modern sequential recommender systems commonly use transformer-based models for next-item prediction. While these models demonstrate a strong balance between efficiency and quality, integrating interleaving features – such as the query context (e.g., browse category) under which next-item interactions occur – poses challenges. Effectively capturing query context is crucial for refining ranking relevance and enhancing user engagement, as it provides valuable signals about user intent within a session. Unlike item features, historical query context is typically not aligned with item sequences and may be unavailable at inference due to privacy constraints or feature store limitations – making its integration into transformers both challenging and error-prone. This paper analyzes different strategies for incorporating query context into transformers trained with a causal language modeling procedure as a case study. We propose a new method that effectively fuses the item sequence with query context within the attention mechanism. Through extensive offline and online experiments on a large-scale online platform and open datasets, we present evidence that our proposed method is an effective approach for integrating query context to improve model ranking quality in terms of relevance and diversity.
@@
Copyright held by the author(s).
copyright: noneconference: Workshop on Context-Aware Recommender Systems in conjunction with RecSys ’25; September 22–26, 2025; Prague, Czech Republicccs: Information systems Recommender systems
@@ 1. Introduction
Existing research on efficient transformer-based sequential models primarily focuses on integrating an item’s features as contextual information (Yang et al., 2021; Cai et al., 2021) or treating query context as side information in conjunction with the user representation, without fusing it early with item sequence (Covington et al., 2016; Zhou et al., 2018), which may limit model effectiveness. When historical query context is assumed to be available online, it has been effectively incorporated into recurrent neural network-based approaches by embedding it within the item sequence, thereby fusing it with the user representation ((Liu et al., 2016; Song et al., 2017; Beutel et al., 2018)).
In this paper, we propose an efficient method to integrate query context into transformer-based sequential models, even when historical context is unavailable at inference. Specifically, we address two research questions:
@@
To answer these, we explore various approaches for incorporating query context into transformer-based sequential recommenders, using causal language modeling as a case study. Our key contributions include a comprehensive study on efficient and effective query context integration, a method that enhances user representation by contextually fusing query context within transformer-based ranking models to improve accuracy, and extensive offline and online experiments on a large-scale platform, validated with open datasets and supported by open-source code for reproducibility.
@@ 2. Related work
Query context is typically incorporated as side information alongside the user representation, rather than being fused early with the item sequence (Covington et al., 2016; Zhou et al., 2018). However, treating query context as side information may limit the predictive power of the user representation, as it is not explicitly tailored to the query context.
@@ 3.1. Problem statement
Here, $N$ is the number of items the user interacted with. During estimation, the $(N+1)$th item is used only as the target when predicting the $N$th next item, while the first item prediction is used for cold-start users.
@@ 3.2. Method
To efficiently model the problem, we consider the deep self-attention transformer model trained with causal language modeling, following the SASRec-style architecture (Kang and McAuley, 2018; Li et al., 2020; Zhang et al., 2019). A transformer begins with an item embedding, followed by a series of self-attention layers. These models stack multiple self-attention layers, each followed by a position-wise feed-forward sublayer, residual connections, and layer normalization (Ba et al., 2016; Nguyen and Salazar, 2019). Multi-head self-attention is commonly used to capture different aspects of input sequences (Vaswani et al., 2017), though we omit it here for notational simplicity.
@@ 4.1. Fixing temporal misalignment
where $\bm{L}$ is a modified lower shift binary NN-by-$(N+1)$ matrix with ones on the subdiagonal. The projections (1) in subsequent layers $h=2,\dots,H$ remain unchanged. This ensures $c_{i+1}$ is correctly placed at query position $\bm{Q}^{(1)}_{i}$, influencing attention over past items $\left(x_{t}^{u}\right)_{t=1}^{i}$ while also propagating the query context further through keys and values. However, this approach raises concerns:
@@ 5.3. Approach C: Query context in the last layer’s query position and outside
Analyzing attention flow in causal language modeling, we conclude that the safest way to introduce query context within attention without causing unintended misrepresentations (Section 4.1) is by placing it in the last layer, exclusively at the query position (Vlasov, 2024).
@@
while the projections in the previous layers, $1=2,\dots,H-1$, remain unchanged as defined in Eq. (1).
@@ 6. Experiments on our large-scale online platform
The model follows a two-tower architecture (Covington et al., 2016; Yang et al., 2020), with the training procedure described in Section 3.2. It consists of a user tower, represented by a transformer with residual layer blocks, and an item tower, represented by a feed-forward network, whose output is the item embedding $\bm{A}$ from Eq. 3. Although these towers are trained jointly, they are deployed and operated independently. The item embeddings are indexed in a vector store for efficient ranking and retrieval. Each time a user makes a request, the user tower generates a new user embedding based on their most recent interactions (up to the last 100 items). This embedding is then used to retrieve the most relevant items using a nearest-neighbor index.
@@ 6.2. Online experiments
    For high-value actions (add-to-wishlist and add-to-cart), both new and returning users saw positive impacts: +1.0% [0.2, 1.8]% for new and +1.7% [1.4 , 2.0]% for returning users. This suggests that the treatment benefits even users for whom we have little to no prior data.
@@ 9. Conclusion
Acknowledgement
We are grateful for the valuable feedback, insightful discussions, and constant support from our many colleagues, as well as their contributions to the design and execution of the online experiments, including Geraud Le Falhe, Danilo Ascione, Tural Gurbanov, Egor Malykh, Gabriel Coelho, Shashi Kumar, Tao Ruangyam, and John Coleman.